# The Multi-Range Theory of Translation Quality Measurement: MQM scoring models and Statistical Quality Control

### v4

@@ 2 On Sampling
For any statistical approach to be applicable, it is critical to know what statistical distribution is valid before choosing the right distribution for further statistical analysis. Such analysis is based on the notion that (a) all errors are independent and (b) the probability of errors in the text is uniform. Although these assumptions are always made, both assumptions are not true concerning texts (and translation products) in general.
@@
There are about a dozen very important factors that can influence why error distribution in text is not, in fact, uniform. Because of this, it is always recommended to revise, review, and evaluate the entire text. But no one has the resources to fully evaluate everything. That’s why translation quality evaluation is typically based on evaluating samples (Gladkoff et al.,, 2022). Sometimes this sampling is done by selecting full-text samples from a larger population, e.g. evaluating every tenth translation fully, or selecting complete shorter texts to arrive at the determined sample size. A typical approach in a traditional localization process is to select samples of medium sizes (500-5000 words) and then apply an analytical approach. However, smaller samples and larger samples are also possible, and it is important to consider the entire range of possible sample sizes, from one sentence to very large documents, which is the purpose of this paper.
@@ 3 Low IRR is not a bug, but a feature
Human translation quality evaluations on small samples have low Inter-Rater Reliability (IRR). What matters in human evaluation is that trained linguists tend to agree (demonstrate high IRR) on bigger samples, but on a segment level two linguists are bound to have three opinions on the same issue. This phenomenon has been widely recognized by both experts in the translation industry and data science (Gladkoff et al., 2023b, ). Research in the field of translation error annotation has demonstrated that linguists disagree even on the precise scope of errors, as well as about error categorizations and severity attributions (Lommel et al., 2014b, ). The researchers established that these issues are not a consequence of low linguist qualifications or technical problems. They are inherent in language itself.
@@
Data scientists often make shallow conclusions from low IRR, believing that automatic calculation of errors will resolve this problem. In recent years, many automatic metrics have been constructed, often claiming “human judgment” quality. Recently, AI-based metrics have appeared, along with unverified and unproven claims that “GenAI seems capable of measuring translation quality” (Gladkoff et al., 2023a, ). In reality, however, GenAI does neither “think” nor “understand” anything and for this reason, the factual accuracy of GenAI generation remains a huge problem. No language model is Turing complete  [^2] https://ai-lab.logrusglobal.com/why-no-agi-can-be-built-with-language-models/. Language models are constrained by their architecture and the limitations of their training data and computational resources. Language models are specialized tools designed for processing and generating text based on patterns learned from vast amounts of data. They are good at some tasks such as language generation, and summarization, but they are not capable of performing arbitrary computations or reasoning in the same way that a Turing complete system can. Because of this, language models are not considered general-purpose reasoners. They can provide useful responses and assist with many tasks within their scope of capabilities, but they cannot reason and compute in the general-purpose manner that a Turing machine or Turing complete system can. Among other things, the factual accuracy of GenAI generation remains to be a huge problem. This affects both error annotation capability and accuracy on a secondary process level. GenAI may produce fluent text, but it performs worse on derivative tasks, such as error annotation tasks. In other words, GenAI misses factual errors when generating content, and of course, does not “see” errors during the annotation process, which also results in issues with the accuracy of this process. Additionally, the better the GenAI output is, the more variable it becomes, which is similar (although different in nature) to the variability in human judgment. Human judgment is variable because different people may have their own interpretation of the text. Advanced GenAI behaves similarly, with its variable output becoming another interpretation. However, this interpretation is not verified or supported by human intelligence.
@@ 4 Three sample-size ranges – three very different methods
The industrial era has developed a vast mathematical and methodological apparatus for measuring the quality of products by evaluating small samples from very large production lots, e.g. using student’s t-distribution by Gladkoff et al., 2023b . In a setting where decisions about the quality of a large lot are based on small samples, the uncertainty is so high that only sophisticated methods of statistical quality control can handle such a problem. They have been extensively developed and described by many researchers, such as Montgomery, (2019), and have been long standardized by ISO  [^3] ISO 2859-2:2020 Sampling procedures for inspection by attributes https://www.iso.org/standard/64505.html. In a nutshell, errors always have a statistical nature, and this is yet another reason why segment-level quality scores do not make sense – not only the methods of producing them must be properly benchmarked and verified, but the fundamental uncertainty of individual error annotation is so high that it is not possible to get a “true” measurement except for very trivial mechanical spelling errors.
@@ Appendix C Highlighted Scorecards and Model Parameters
Figure 9: Quality Measure and Tools (red squares with straight corners), and Quality Scores (orange squares with curved corners).

### v5

@@ 2 On Sampling
For any statistical approach to be applicable, it is critical to know what statistical distribution is valid before choosing the right distribution for further statistical analysis. Such analysis is based on the notion that, ideally, (a) all errors are independent and (b) the probability of errors in the text is uniform. Although these assumptions are always made, neither assumption is true concerning texts (and translation products) in general.
@@
There are about a dozen very important factors that can influence why error distribution in text is not, in fact, uniform. Because of this, it is always recommended to revise, review, and evaluate the entire text. But no one has the resources to fully evaluate everything. That’s why translation quality evaluation is typically based on evaluating samples (such as the work by Gladkoff et al., (2022) from an industrial setting). Sometimes this sampling is done by selecting full-text samples from a larger population, e.g. evaluating every tenth translation fully, or selecting complete shorter texts to arrive at the determined sample size. A typical approach in a traditional localization process is to select samples of medium sizes (500-5000 words) and then apply an analytical approach. However, smaller samples and larger samples are also possible, and it is important to consider the entire range of possible sample sizes, from one sentence to very large documents, which is the purpose of this paper.
@@ 3 Low IRR is not a bug, but a feature
Human translation quality evaluations on small samples have low Inter-Rater Reliability (IRR). What matters in human evaluation is that trained linguists tend to agree (demonstrate high IRR) on bigger samples, but on a segment level two linguists often disagree on the same issue. This phenomenon has been widely recognized by both experts in the translation industry and data science (Gladkoff et al.,, 2023). Research in the field of translation error annotation has demonstrated that linguists can disagree even on the precise scope of errors, as well as about error categorizations and severity attributions (Lommel et al., 2014b, ). The researchers established that these issues are not a consequence of low linguist qualifications or technical problems. Multiple factors at play can determine disagreement among annotators: 1) the complexity of the task; 2) the ambiguity of the text; 3) the quality of the translation; 4) subjectivity; etc.
@@
Data scientists often make shallow conclusions from low IRR, believing that automatic calculation of errors will resolve this problem. In recent years, many automatic metrics have been constructed, often claiming “human judgment” quality. Recently, AI-based metrics have appeared, along with unverified and unproven claims that “GenAI seems capable of measuring translation quality” (Gladkoff et al.,, 2024). In reality, however, GenAI does neither “think” nor “understand” anything and for this reason, the factual accuracy of GenAI generation remains a huge problem. No language model is Turing complete  [^2] https://ai-lab.logrusglobal.com/why-no-agi-can-be-built-with-language-models/. Language models are constrained by their architecture and the limitations of their training data and computational resources. Language models are specialized tools designed for processing and generating text based on patterns learned from vast amounts of data. They are good at some tasks such as language generation, and summarization, but they are not capable of performing arbitrary computations or reasoning in the same way that a Turing complete system can. Because of this, language models are not considered general-purpose reasoners. They can provide useful responses and assist with many tasks within their scope of capabilities, but they cannot reason and compute in the general-purpose manner that a Turing machine or Turing complete system can. Among other things, the factual accuracy of GenAI generation remains to be a huge problem. This affects both error annotation capability and accuracy on a secondary process level. GenAI may produce fluent text, but it performs worse on derivative tasks, such as error annotation tasks. In other words, GenAI misses factual errors when generating content, and of course, does not “see” errors during the annotation process, which also results in issues with the accuracy of this process. Additionally, the better the GenAI output is, the more variable it becomes, which is similar (although different in nature) to the variability in human judgment. Human judgment is variable because different people may have their own interpretation of the text. Advanced GenAI behaves similarly, with its variable output becoming another interpretation. However, this interpretation is not verified or supported by human intelligence.
@@ 4 Three sample-size ranges – three very different methods
The industrial era has developed a vast mathematical and methodological apparatus for measuring the quality of products by evaluating small samples from very large production lots, e.g. using student’s t-distribution (Student,, 1908) such as the recent work by Gladkoff et al., (2023). In a setting where decisions about the quality of a large lot are based on small samples, the uncertainty is so high that only sophisticated methods of statistical quality control can handle such a problem. They have been extensively developed and described by many researchers, such as Montgomery, (2019), and have been long standardized by ISO  [^3] ISO 2859-2:2020 Sampling procedures for inspection by attributes https://www.iso.org/standard/64505.html. In a nutshell, errors always have a statistical nature, and this is yet another reason why segment-level quality scores do not make sense – not only the methods of producing them must be properly benchmarked and verified, but the fundamental uncertainty of individual error annotation is so high that it is not possible to get a “true” measurement except for very trivial mechanical spelling errors.
@@ 5.2 Scoring Model
Figure 1: MQM2.0 Evaluation Scorecard: Quality Measure and Tools (red squares with straight corners), and Quality Scores (orange squares with curved corners).
@@ 8 Conclusion and Future Work
Acknowledgements
We thank the reviewers for their valuable comments and feedback on our work. LH and GN are grateful for the grant “Integrating hospital outpatient letters into the healthcare data space” (EP/V047949/1; funder: UKRI/EPSRC).